# Large Stepsize Gradient Descent for Non-Homogeneous Two-Layer Networks: Margin Improvement and Fast Optimization

### v1

@@ 1 Introduction
Figure 1: The behavior of (GD) for optimizing a non-homogenous MLP with tanh activation function on a subset of CIFAR-10 dataset. We randomly sample $1,979$ data with labels “airplane” and “automobile” from CIFAR-10 dataset. The normalized margin is defined by 3. The blue curves correspond to GD with a large stepsize $\tilde{\eta}=1$, where the empirical risk oscillates in the first phase but decreases monotonically in the second phase. The orange curves correspond to GD with a small stepsize $\tilde{\eta}=0.02$, where the empirical risk decreases monotonically. Furthermore, Figure 1(b) suggests the normalized margins of both two curves increase and converge in the stable phases. Finally, Figure 1(c) suggests that large stepsize achieves a better test accuracy, consistent with larger-scale learning experiment [Goyal et al., 2017]. More details can be found in Section 5.
@@
Here, the predictor $f(\mathbf{w};\cdot):\mathbb{R}^{d}\mapsto\mathbb{R}$ is parameterized by trainable parameters $\mathbf{w}$ and is assumed to be continuously differentiable with respect to $\mathbf{w}$. The predictor is initialized from $\mathbf{w}_{0}$ and then trained by gradient descent (GD) with a constant stepsize $\tilde{\eta}>0$, that is,
@@
    Assuming that GD enters the stable phase, if in addition the preditor has a bounded homogenous error (see 1C), we show that the normalized margin induced by GD, $\min_{i}y_{i}f(\mathbf{w}_{t};\mathbf{x}_{i})/\|\mathbf{w}_{t}\|$, nearly monotonically increases (see Theorem 2.2). To the best of our knowledge, this is the first characterization of implicit bias of GD for non-homogenous predictors. In particular, our theory covers two-layer networks with commonly used activations functions (which are often non-homogenous) that cannot be covered by existing results [Lyu and Li, 2020, Ji and Telgarsky, 2020, Chizat and Bach, 2020].
@@ 2 Stable Phase and Margin Improvement
    Near homogeneity. Assume there exists $\kappa>0$ such that for every $\mathbf{w}$,
@@
  • Sigmoid. $\phi(x):={1}/{(1+e^{-x})}$ with $\kappa=1$, $\beta=2/m$, and $\rho=1/\sqrt{m}$.
  • Tanh. $\phi(x):=\tanh(x)$ with $\kappa=5$, $\beta=2/m$, and $\rho=1/\sqrt{m}$.
@@
  • Risk convergence. $L(\mathbf{w}_{t})=\Theta(1/t)$, where the constants depend on $L(\mathbf{w}_{s}),\|\mathbf{w}_{s}\|,\tilde{\eta}$ and $\rho$.
  • Parameter increase. $\|\mathbf{w}_{t+1}\|\geq\|\mathbf{w}_{t}\|$ and $\|\mathbf{w}_{t}\|=\Theta(\log(t))$, where the constants depend on $n,\rho,\kappa,\tilde{\eta},\|\mathbf{w}_{s}\|,L(\mathbf{w}_{s})$ and $f(\mathbf{w}_{s};\mathbf{x}_{i})$.
@@
Theorem 2.2 shows that for an arbitrarily large stepsize $\tilde{\eta}$, GD must enter the stable phase if the empirical risk falls below a threshold depending on $\tilde{\eta}$ given by 4. In the next section, We will show that under additional technical conditions, large stepsize GD is guaranteed to satisfy 4 even though with an initially oscillatory risk.
Furthermore, for nearly homogenous predictors, Theorem 2.2 shows that under a stronger risk threshold condition 5, the risk must converge at a $\Theta(1/t)$ rate and that the normalized margin nearly monotonically increases. This demonstrates an implicit bias of GD, even when used with a large stepsize and the trained predictor is non-homogenous.
@@ 3 Edge of Stability Phase
    Near homogeneity. Assume there exists $\kappa>0$ such that for every $z\in\mathbb{R}$, $|\phi(z)-\phi^{\prime}(z)z|\leq\kappa$.
@@
Fix $10.5\leq c<$ 1. For each $\phi$ in Example 2.1 except the Huberized ReLU, its modification $\tilde{\phi}(x):=cx+(1-c)/4\cdot\phi(x)$ satisfies Assumption 2 with $\kappa=1$, $\alpha=0.25$, and $\tilde{\beta}=1$. The modification of the Huberized ReLU satisfies Assumption 2 with $\kappa=h/2$, $\alpha=0.5$, and $\tilde{\beta}=1/4h$. In particular, the modification of softplus can be viewed as a smoothed leaky ReLU.
@@ 5 Experiments
We conduct three sets of experiments to validate our theoretical insights. In the first set, we use a subset of the CIFAR-10 dataset [Krizhevsky et al., 2009], which includes 1,979 randomly selected samples from the ”airplane” and ”automobile” classes. Our model is a multilayer perceptron (MLP) with two trainable layers and tanh activation functions, with a hidden dimension of 100. The MLP is trained using gradient descent with random initialization, as described in GD. The results are shown in Figures 1(a), 1(c) and 1(b).
@@ 6 Related Works
For logistic regression on linearly separable data, Soudry et al. [2018], Ji and Telgarsky [2018] showed that the direction of small stepsize GD converges to the max-margin solution. Their results were later extended by Gunasekar et al. [2017, 2018], Nacson et al. [2019c, a, b], Ji et al. [2021], Lyu and Li [2020], Ji and Telgarsky [2020], Chizat and Bach [2020] to other algorithms and non-linear models. However, in all of their analysis, the stepsize of GD needs to be small such that the empirical risk decreases monotonically. In contrast, our focus is GD with a large stepsize that induces non-monotonic risk.
@@
In practice, large stepsizes are often preferred when using GD to train neural networks to achieve effective optimization and generalization performance [see Wu and Ma, 2018, Cohen et al., 2020, and references therein]. In such scenarios, the empirical risk often oscillates in the beginning. This phenomenon is named edge of stability (EoS) by Cohen et al. [2020]. The theory of EoS is mainly studied in relatively simplified cases such as one- or two-dimensional functions [Zhu et al., 2022, Chen and Bruna, 2023, Ahn et al., 2022, Kreisler et al., 2023, Wang et al., 2023], linear model [Wu et al., 2023, 2024], matrix factorization [Wang et al., 2022a, Chen and Bruna, 2023], scale-invariant networks [Lyu et al., 2022], for an incomplete list of references. Compared to them, we focus on a more practical setup of training two-layer non-linear networks with large stepsize GD. There are some general theories of EoS subject to subtle assumptions [Kong and Tao, 2020, Ahn et al., 2022, Ma et al., 2022, Damian et al., 2022, Wang et al., 2022b], which are not directly comparable to ours.
@@
Both results in [Lyu and Li, 2020, Ji and Telgarsky, 2020] focused on $L$-homogenous networks. Specifically, Lyu and Li [2020] showed that a modified version of normalized margin (see 3) induced by GD with small stepsize (such that the risk decreases monotonically) increases, with limiting points of $\{\mathbf{w}_{t}/\|\mathbf{w}_{t}\|\}_{t=1}^{\infty}$ converging to KKT points of a margin-maximization problem. Under additional o-minimal conditions, Ji and Telgarsky [2020] showed that gradient flow converges in direction. Our work is different from theirs in two aspects. First, we allow GD with a large stepsize that may cause risk oscillation. Second, our theory covers non-homogenous predictors, which include two-layer networks with many commonly used activation functions beyond the scope of [Lyu and Li, 2020, Ji and Telgarsky, 2020]. Compared to Lyu and Li [2020], Ji and Telgarsky [2020], we only show the improvement of the margin, and our theory is limited to nearly $1$-homogenous predictors (2C). It remains open to show directional convergence and to extend our near $1$-homogenity condition to a “near $L$-homogeneity” condition.
@@
The work by Chatterji et al. [2021] studies the convergence of GD in training deep networks under logistic loss. Their results are related to ours as we both consider networks with nearly homogeneous activations and we both have a stable phase analysis (although this is not explicitly mentioned in their paper). However, our results are significantly different from theirs. Specifically, in our notation, they require the homogenous error $\kappa$ (see 2C) to be smaller than $\mathcal{O}(\log(1/L(\mathbf{w}_{s}))/\|\mathbf{w}_{s}\|)\approx\mathcal{O}(\bar{\gamma}(\mathbf{w}_{s}))$, where $s$ is the time for GD to enter the stable phase. Note that the margin when GD enters the stable phase could be arbitrarily small. In comparison, we only require the homogenous error to be bounded by a constant. As a consequence, we can handle many commonly used activation functions (see Example 2.1) while they can only handle the Huberized ReLU with a small $h$ in Example 2.1. Moreover, they require the stepsize $\tilde{\eta}$ to be smaller than $\mathcal{O}(\kappa/\|\mathbf{w}_{s}\|^{8})$, thus they only allow very small stepsize. In contrast, we allow $\tilde{\eta}$ to be an arbitrarily large constant.
@@ Acknowledgements
We thank Fabian Pedregosa for his suggestions on an early draft. We gratefully acknowledge the support of the NSF for FODSI through grant DMS-2023505, of the NSF and the Simons Foundation for the Collaboration on the Theoretical Foundations of Deep Learning through awards DMS-2031883 and #814639, and of the ONR through MURI award N000142112431.

### v2

@@ 1 Introduction
Figure 1: The behavior of (GD) for optimizing a non-homogenous four-layer MLP with GELU activation function on a subset of CIFAR-10 dataset. We randomly sample $6,000$ data with labels “airplane” and “automobile” from CIFAR-10 dataset. The normalized margin is defined as $\big{(}\arg\min_{i\in[n]}y_{i}f(\mathbf{w}_{t};\mathbf{x}_{i})\big{)}/\|\mathbf{w}_{t}\|^{4}$, which is close to 3. The blue curves correspond to GD with a large stepsize $\tilde{\eta}=0.2$, where the empirical risk oscillates in the first phase but decreases monotonically in the second phase. The orange curves correspond to GD with a small stepsize $\tilde{\eta}=0.005$, where the empirical risk decreases monotonically. Furthermore, Figure 1(b) suggests the normalized margins of both two curves increase and converge in the stable phases. Finally, Figure 1(c) suggests that large stepsize achieves a better test accuracy, consistent with larger-scale learning experiment [Hoffer et al., 2017, Goyal et al., 2017]. More details can be found in Section 5.
@@
Here, the predictor $f(\mathbf{w};\cdot):\mathbb{R}^{d}\mapsto\mathbb{R}$ is parameterized by trainable parameters $\mathbf{w}$ and is assumed to be continuously differentiable for $\mathbf{w}$. The predictor is initialized from $\mathbf{w}_{0}$ and then trained by gradient descent (GD) with a constant stepsize $\tilde{\eta}>0$, that is,
@@
    Assuming that GD enters the stable phase, if in addition the preditor has a bounded homogenous error (see 1C), we show that the normalized margin induced by GD, $\min_{i}y_{i}f(\mathbf{w}_{t};\mathbf{x}_{i})/\|\mathbf{w}_{t}\|$, nearly monotonically increases (see Theorem 2.2). This characterizes an implicit bias of GD for non-homogenous predictors. In particular, our theory covers two-layer networks with non-homogenous activations such as GeLU and SiLU that cannot be covered by existing results [Lyu and Li, 2020, Ji and Telgarsky, 2020, Chizat and Bach, 2020].
@@ 2 Stable Phase and Margin Improvement
    Near-homogeneity. Assume there exists $\kappa>0$ such that for every $\mathbf{w}$,
@@
  • Risk convergence. $L(\mathbf{w}_{t})=\Theta(1/t)$, where $\Theta$ hides the dependence on $L(\mathbf{w}_{s}),\|\mathbf{w}_{s}\|,\tilde{\eta}$ and $\rho$.
  • Parameter increase. $\|\mathbf{w}_{t+1}\|\geq\|\mathbf{w}_{t}\|$ and $\|\mathbf{w}_{t}\|=\Theta(\log(t))$, where $\Theta$ hides the dependence on $n$, $\rho$, $\kappa$, $\tilde{\eta}$, $\|\mathbf{w}_{s}\|$, $L(\mathbf{w}_{s})$ and $f(\mathbf{w}_{s};\mathbf{x}_{i})$ for $i=1,\dots,n$.
@@
Theorem 2.2 shows that for an arbitrarily large stepsize $\tilde{\eta}$, GD must enter the stable phase if the empirical risk falls below a threshold depending on $\tilde{\eta}$ given by 4. Furthermore, for nearly homogenous predictors, Theorem 2.2 shows that under a stronger risk threshold condition 5, the risk must converge at a $\Theta(1/t)$ rate and that the normalized margin nearly monotonically increases. This demonstrates an implicit bias of GD, even when used with a large stepsize and the trained predictor is non-homogenous.
Limitations.
The stable phase conditions in Theorem 2.2 require GD to enter a sublevel set of the empirical risk. However, such a sublevel set might be empty. For instance, let $f(\mathbf{w};\mathbf{x})$ be a two-layer network 2 with sigmoid activation. Notice that the predictor is uniformly bounded, $|f(\mathbf{w};\mathbf{x})|\leq 1$, so we have
$L(\mathbf{w})=\frac{1}{n}\sum_{i=1}^{n}\log(1+e^{-y_{i}f(\mathbf{w};\mathbf{x}_{i})})\geq\log(1+e^{-1}).$
On the other hand, we can also verify that 1C is satisfied by $f(\mathbf{w};\mathbf{x})$ with $\kappa=1$ but no smaller $\kappa$. Therefore 5 cannot be satisfied. In general, the sublevel set given by the right-hand side of 1C is non-empty if
[…]
The above condition requires the data can be separated arbitrarily well by some predictor within the hypothesis class. This condition is general and covers (sufficiently large) two-layer networks 2 with many commonly used activations such as GeLU and SiLU. Moreover, although two-layer networks with sigmoid activation violate this condition, they can be modified by adding a leakage to the sigmoid to satisfy the condition.
In the next section, we will provide sufficient conditions such that large stepsize GD will enter the stable phases characterized by 4 or 5.
Comparisons to existing works.
@@ 3 Edge of Stability Phase
    Near-homogeneity. Assume there exists $\kappa>0$ such that for every $z\in\mathbb{R}$, $|\phi(z)-\phi^{\prime}(z)z|\leq\kappa$.
@@
Fix $10.5\leq c<$ 1.
  • Let $\phi$ be GELU, Softplus, or SilU in Example 2.1, then its modification $\tilde{\phi}(x):=cx+(1-c)/4\cdot\phi(x)$ satisfies Assumption 2 with $\kappa=1$, $\alpha=0.25$, and $\tilde{\beta}=1$. In particular, the modification of softplus can be viewed as a smoothed leaky ReLU.
  • Let $\phi$ be the Huberized ReLU in Example 2.1, then its modification $\tilde{\phi}(x):=cx+(1-c)/4\cdot\phi(x)$ satisfies Assumption 2 with $\kappa=h/2$, $\alpha=0.5$, and $\tilde{\beta}=1/4h$.
  • The “leaky” tanh, $\tilde{\phi}(x)\coloneqq cx+(1-c)\tanh(x)$, and the “leaky” sigmoid, $\tilde{\phi}(x)\coloneqq cx+{c}/{(1+e^{-x})}$, both satisfy Assumption 2 with $\kappa=1,\alpha=0.5$ and $\tilde{\beta}=1$.
@@ 5 Experiments
We conduct three sets of experiments to validate our theoretical insights. In the first set, we use a subset of the CIFAR-10 dataset [Krizhevsky et al., 2009], which includes 6,000 randomly selected samples from the “airplane” and “automobile” classes. Our model is a multilayer perceptron (MLP) with four trainable layers and GELU activation functions, with a hidden dimension of 200 for each hidden layer. The MLP is trained using gradient descent with random initialization, as described in GD. The results are shown in Figures 1(a), 1(c) and 1(b).
@@ 6 Related Works
For logistic regression on linearly separable data, Soudry et al. [2018], Ji and Telgarsky [2018] showed that the direction of small stepsize GD converges to the max-margin solution. Their results were later extended by Gunasekar et al. [2017, 2018], Nacson et al. [2019c, a, b], Ji et al. [2021], Lyu and Li [2020], Ji and Telgarsky [2020], Chizat and Bach [2020], Chatterji et al. [2021], Kunin et al. [2022] to other algorithms and non-linear models. However, in all of their analysis, the stepsize of GD needs to be small such that the empirical risk decreases monotonically. In contrast, our focus is GD with a large stepsize that induces non-monotonic risk.
Two papers [Nacson et al., 2019a, Kunin et al., 2022] studied margin maximization theory for a special form of non-homogenous models. Specifically, when viewed in terms of different subsets of the trainable parameters, the model is homogeneous, although the order of homogeneity may vary. Compared to their setting, our non-homogenous models only require a bounded homogenous error (see 1C). Therefore, our theory can cover two-layer networks 2 with non-homogeneous activations such as GELU and SiLU that cannot be covered by [Nacson et al., 2019a, Kunin et al., 2022].
@@
In practice, large stepsizes are often preferred when using GD to train neural networks to achieve effective optimization and generalization performance [see Wu and Ma, 2018, Cohen et al., 2020, and references therein]. In such scenarios, the empirical risk often oscillates in the beginning. This phenomenon is named edge of stability (EoS) by Cohen et al. [2020]. The theory of EoS is mainly studied in relatively simplified cases such as one- or two-dimensional functions [Zhu et al., 2022, Chen and Bruna, 2023, Ahn et al., 2022, Kreisler et al., 2023, Wang et al., 2023], linear model [Wu et al., 2023, 2024], matrix factorization [Wang et al., 2022a, Chen and Bruna, 2023], scale-invariant networks [Lyu et al., 2022], for an incomplete list of references. Compared to them, we focus on a more practical setup of training two-layer non-linear networks with large stepsize GD. There are some general theories of EoS subject to subtle assumptions [for example, Kong and Tao, 2020, Ahn et al., 2022, Ma et al., 2022, Damian et al., 2022, Wang et al., 2022b, Lu et al., 2023], which are not directly comparable to ours.
@@
Both results in [Lyu and Li, 2020, Ji and Telgarsky, 2020] focused on $L$-homogenous networks. Specifically, Lyu and Li [2020] showed that a modified version of normalized margin (see 3) induced by GD with small stepsize (such that the risk decreases monotonically) increases, with limiting points of $(\mathbf{w}_{t}/\|\mathbf{w}_{t}\|)_{t=1}^{\infty}$ converging to KKT points of a margin-maximization problem. Under additional o-minimal conditions, Ji and Telgarsky [2020] showed that gradient flow converges in direction. Our work is different from theirs in two aspects. First, we allow GD with a large stepsize that may cause risk oscillation. Second, our theory covers non-homogenous predictors, which include two-layer networks with many commonly used activation functions beyond the scope of [Lyu and Li, 2020, Ji and Telgarsky, 2020]. Compared to Lyu and Li [2020], Ji and Telgarsky [2020], we only show the improvement of the margin, and our theory is limited to nearly $1$-homogenous predictors (2C). It remains open to show directional convergence and to extend our near $1$-homogenity condition to a “near $L$-homogeneity” condition.
@@
The work by Chatterji et al. [2021] studied the convergence of GD in training deep networks under logistic loss. Their results are related to ours as we both consider networks with nearly homogeneous activations and we both have a stable phase analysis (although this is not explicitly mentioned in their paper). However, our results are significantly different from theirs. Specifically, in our notation, they require the homogenous error $\kappa$ (see 2C) to be smaller than $\mathcal{O}(\log(1/L(\mathbf{w}_{s}))/\|\mathbf{w}_{s}\|)\approx\mathcal{O}(\bar{\gamma}(\mathbf{w}_{s}))$, where $s$ is the time for GD to enter the stable phase. Note that the margin when GD enters the stable phase could be arbitrarily small. In comparison, we only require the homogenous error to be bounded by a constant. As a consequence, we can handle many commonly used activation functions (see Example 2.1) while they can only handle the Huberized ReLU with a small $h$ in Example 2.1. Moreover, they require the stepsize $\tilde{\eta}$ to be smaller than $\mathcal{O}(\kappa/\|\mathbf{w}_{s}\|^{8})$, thus they only allow very small stepsize. In contrast, we allow $\tilde{\eta}$ to be an arbitrarily large constant.
@@ Acknowledgements
We thank Fabian Pedregosa for his suggestions on an early draft and Jason D. Lee and Kaifeng Lyu for their comments clarifying the applicability of our result to sigmoid networks. We gratefully acknowledge the support of the NSF for FODSI through grant DMS-2023505, of the NSF and the Simons Foundation for the Collaboration on the Theoretical Foundations of Deep Learning through awards DMS-2031883 and #814639, and of the ONR through MURI award N000142112431.